4. If a random sample of 50 cups of soda dispensed from a machine showed the mean amount of soda dispensed was 7.9 ounces with s=0.35 ounces, find the confidence level if the margin of error is 0.083.
5. Suppose that a random sample of 25 mining caps was tested from a population of 300 mining caps and 20 exploded properly.
1
Expert's answer
2020-10-05T18:10:54-0400

CI =X̄ ± Z×s/"\sqrt n" =7.9 ± 0.083

Z×s/"\sqrt n" =Z×( 0.35"/\sqrt {50}" )=0.083

Z=1.64, using table, confidence level=90%